OpenStudy (anonymous):

what is the domain of x : x^{1-\frac{1}{3}x^2} = \frac{1}{\sqrt[3]{100}}

OpenStudy (anonymous):

could clarify the question.

OpenStudy (anonymous):

\[\huge x^{1-\frac{1}{3}x^2} = \frac{1}{\sqrt[3]{100}}\]

OpenStudy (anonymous):

yes